# Advancing Graph Generation through Beta Diffusion

### v1

@@ Results.
For Ego-small and Community-small, it is worth noting that for most evaluation metrics, the MMDs of the graphs sampled from the training set still have large standard deviations, likely due to the small number of nodes in the graph and the insufficient size of the sampled graphs. Therefore, it is necessary to assess the quality of generated samples more definitively using a larger number of generated graphs compared to the test data. Similar to the previous works (Cho et al., 2023; Jo et al., 2022; Liu et al., 2019), we sampled 1024 graphs for each smaller dataset and evaluated the MMD metrics with their means and standard deviations reported in Table 2. We observe that our proposed GBD outperforms previous continuous and discrete diffusion models on both smaller datasets. Furthermore, GBD significantly surpasses the wavelet-based diffusion model (Wave-GD) by a wide margin on the Community-small dataset, as evidenced by both means and standard deviations. Specifically, GBD achieves 85.0%, 90.5%, and 40.0% improvements over Wave-GD in the MMDs means of Degree, Cluster, and Orbit, respectively, indicating that our model is capable of generating smaller graphs that are closer to the data distribution with better stability.
Table 1: Generic graph generation results.
Ego-small Community-small Grid
Method Deg. $\downarrow$ Clus. $\downarrow$ Orbit. $\downarrow$ Avg. $\downarrow$ Deg. $\downarrow$ Clus. $\downarrow$ Orbit. $\downarrow$ Avg. $\downarrow$ Deg. $\downarrow$ Clus. $\downarrow$ Orbit. $\downarrow$ Avg. $\downarrow$
DeepGMG 0.040 0.100 0.020 0.053 0.220 0.950 0.400 0.523 - - - -
GraphRNN 0.090 0.220 0.003 0.104 0.080 0.120 0.040 0.080 0.064 0.043 0.021 0.043
GraphAF 0.030 0.110 0.001 0.047 0.180 0.200 0.020 0.133 - - - -
GraphDF 0.040 0.130 0.010 0.060 0.060 0.120 0.030 0.070 - - - -
GraphVAE 0.130 0.170 0.050 0.117 0.350 0.980 0.540 0.623 1.619 0.0 0.919 0.846
GNF 0.030 0.100 0.001 0.044 0.200 0.200 0.110 0.170 - - - -
EDP-GNN 0.052 0.093 0.007 0.051 0.053 0.144 0.026 0.074 0.455 0.238 0.328 0.340
GDSS 0.021 0.024 0.007 0.017 0.045 0.086 0.007 0.046 0.111 0.005 0.070 0.062
ConGress 0.037 0.064 0.017 0.039 0.020 0.076 0.006 0.034
DiGress 0.017 0.021 0.010 0.016 0.028 0.115 0.009 0.050 - - - -
Wave-GD 0.012 0.010 0.005 0.009 0.007 0.058 0.002 0.022 0.144 0.004 0.021 0.056
GBD 0.011 0.014 0.002 0.009 0.002 0.060 0.002 0.021 0.045 0.011 0.040 0.032
($\pm$0.008) ($\pm$0.013) ($\pm$0.004) - ($\pm$0.003) ($\pm$0.004) ($\pm$0.003) - ($\pm$0.004) ($\pm$0.002) ($\pm$0.014) -
@@
Table 2: Generic graph generation results with larger sample size.
@@
4.2 Molecule Generation
We validate GBD on 2D molecule generation tasks for attributed graphs, demonstrating its capability to model graph structure with both node and edge attributes.
Datasets and Metrics
Following the evaluation setting of Jo et al. (2022), we generated 10,000 molecules for each dataset and evaluate them with four metrics: the ratio of valid molecules without correction (Val.), Fréchet ChemNet Distance (FCD), Neighborhood subgraph pairwise distance kernel NSPDK, and Scaffold similarity (Scaf.).
Baselines
We compare GBD against the following autoregressive and one-shot graph generation methods: MoFlow (Zang and Wang, 2020), GraphAF (Shi et al., 2020), GraphDF (Luo et al., 2021), and several state-of-the-art diffusion-based graph generative models discussed previously: EDP-GNN (Niu et al., 2020a), GDSS (Jo et al., 2022) and ConGress (Vignac et al., 2023), DiGress (Vignac et al., 2023), and DruM (Jo et al., 2023). We describe the implementation details in Appendix A.
@@ Results
As shown in Table 3, we observe that our GBD outperforms most previous diffusion-based models and is competitive with the current state-of-the-art Gaussian-based diffusion model, DruM. In particular, compared to the basic continuous diffusion model, GBD significantly outperforms it (GDSS+Transformer) under the same GraphTransformer architecture. Additionally, we observe that our proposed beta-based diffusion model is superior to the discrete diffusion model on both 2D molecule datasets, demonstrating that the beta-based diffusion model is also capable of modeling complex structures of attributed graphs (even more effectively when comparing GBD to DiGress). We attribute this to the excellent modeling ability of beta-based diffusion model for sparse and long-tailed data distributions.
Table 3: 2D molecule generation results
QM9 ($|V|\leq 9$) ZINC250K
Method Valid (%) $\uparrow$ FCD $\downarrow$ NSPDK $\downarrow$ Scaf. $\downarrow$ Valid(%) $\uparrow$ FCD $\downarrow$ NSPDK $\downarrow$ Scaf. $\downarrow$
MoFlow 91.36 4.467 0.0169 0.1447 63.11 20.931 0.0455 0.0133
GraphAF 74.43 5.625 0.0207 0.3046 68.47 16.023 0.0442 0.0672
GraphDF 93.88 10.928 0.0636 0.0978 90.61 33.546 0.1770 0.0000
EDP-GNN 47.52 2.680 0.0046 0.3270 82.97 16.737 0.0485
GDSS 95.72 2.900 0.0033 0.6983 97.01 14.656 0.0195 0.0467
GDSS+Transformer 99.68 0.737 0.0024 0.9129 96.04 5.556 0.0326 0.3205
DiGress 98.19 0.095 0.0003 0.9353 94.99 3.482 0.0021 0.4163
DruM 99.69 0.108 0.0002 0.9449 98.65 2.257 0.0015 0.5299
GBD 99.88 0.093 0.0002 0.9510 97.87 2.248 0.0018 0.5042
4.3 Ablation studies
With all other hyperparameter choices kept constant, we vary the options regarding computation domain and the application of preconditioning, and summarize the outcome of model performance in Table 4. The combination of adopting logit domain computation without using preconditioning can sometimes increase the challenge in model convergence, and therefore, it is not recommended. The listed results on Ego-small and Community-small demonstrate that both techniques are in general beneficial for achieving better model performance, and the effect of preconditioning is more evident when the computation is perfomed in the logit domain.
Table 4: The effect of practicing logit domain computation and preconditioning
Ego-small Community-small
logit domain preconditioning Deg. $\downarrow$ Clus. $\downarrow$ Orbit. $\downarrow$ Avg. $\downarrow$ Deg. $\downarrow$ Clus. $\downarrow$ Orbit. $\downarrow$ Avg. $\downarrow$
- - 0.015 0.018 0.004 0.012 0.010 0.076 0.004 0.030
- 0.013 0.017 0.002 0.011 0.004 0.044 0.007 0.018
0.011 0.014 0.002 0.009 0.002 0.060 0.002 0.021
4.4 Visualization
@@
We provide the visualization of the generative process and generated graphs of GBD in Appendix B. For generative process with general datasets shown in Appendix B.1, we follow the implementation described in Section 2.4 and the nodes in all adjacency matrices are reordered by decreasing node degree. Apparently, we can find that edges associated with nodes with large degree will be the first to be identified, and then spread in decreasing order of degree on both datasets. It’s worth noting that the reverse beta diffusion can converge rapidly, leading to generated graphs with correct topology at an early stage. This demonstrates that our proposed GBD can further explore the potential benefits of beta diffusion, resulting in valid graphs with stability and high quality. For generated molecule graphs shown in Appendix B.1, we can observe that GBD can successfully generate valid and high-quality 2D molecules, verifying its ability to model attributed graphs.
5 Conclusion, Limitations and Broader Impact
We introduce graph beta diffusion (GBD), a novel graph generation framework developed upon beta diffusion. We demonstrate that the utilization of beta distribution to define the diffusion process is beneficial for modeling the distribution of graph data, and outline four crucial designing elements—data transformation, concentration modulation, logit-domain computation, and neural-network precondition—that that consistently enhance model performance.
With these contributions achieved, we identify several areas for potential improvement. First, while the beta distribution can adeptly model various data distributions within the range of $[0,1]$, using it to construct the diffusion model necessitates careful selection of scaling and shifting parameters to ensure model convergence. This requirement can complicate the transfer of experience between different tasks. Second, unlike discrete diffusion processes, the beta diffusion process does not naturally generate discrete intermediate adjacency matrices for computing graph statistics. This necessitates a quantization strategy if one aims to incorporate statistics from these intermediates when predicting the graph at the initial state. Thirdly, diffusion-based graph generative models, including GBD, currently rely on a reverse diffusion process that takes hundreds to thousands of iterative refinement steps to generate a single sample. Recent advancements in score-based distillation techniques, originally developed for image diffusion models, could be adapted to distill the graph teacher model (Luo et al., 2023; Zhou et al., 2024). Such an adaptation has the potential to significantly accelerate the graph generation process while maintaining or even improving performance. We plan to explore these possibilities in future studies.
Finally, we explore the potential impact of GBD. Since much real-world data can be structured as graphs, an effective tool for generating realistic graphs could significantly benefit researchers in the natural and social sciences, enabling the economical creation of high-quality simulated data. However, there is a concern: if the generated content is misused for fraudulent activities, distinguishing it from authentic data could become increasingly challenging for recipients.
Appendix A Experimental Details
A.1 General graph generation
Datasets
We evaluated our model using three synthetic and real datasets of varying size and connectivity, previously used as benchmarks in the literature [Cho et al., 2023, Jo et al., 2022]: Ego-small [Sen et al., 2008] consists of 200 small real sub-graphs from the Citeseer network dataset with $4\leq N\leq 18$. Community-small consists of 100 randomly generated synthetic graphs with $12\leq N\leq 20$, where the graphs are constructed by two equal-sized communities, each of which is generated by the Erdös–Rényi model [Erdős et al., 1960], with $p=0.7$ and $0.05N$ inter-community edges are added with uniform probability as in previous works [Jo et al., 2022, Niu et al., 2020b]. Grid consists of randomly generated 100 standard 2D grid graphs with $100\leq N\leq 400$ and the maximum number of edges per node is 4 since all nodes are arranged in a regular lattice.
Evaluation metrics
For a fair comparison, we follow the experimental and evaluation settings of Jo et al. [2022, 2023], using the same train/test split, where 80% of the data is used as the training set and the remaining 20% as the test set. We adopt maximum mean discrepancy (MMD) as our evaluation metric to compare three graph property distributions between test graphs and the same number of generated graphs: degree (Deg.), clustering coefficient (Clus.), count of orbits with 4 nodes (Orbit) and their average score (Avg.). Note that we use the Gaussian Earth Mover’s Distance (EMD) kernel to compute the MMDs following the method used in previous work [Jo et al., 2022, Cho et al., 2023].
Implementation details
We follow the evaluation setting of  Jo et al. [2022], Cho et al. [2023] to generate graphs of the same size as the test data in each run and we report the mean and standard deviation obtained from 3 independent runs for each dataset. We report the baseline results taken from  Cho et al. [2023], except for the results of ConGress in Tables 1 and 2, which we obtained by running its corresponding open-source code. For a fair comparison, we adopt the Graph Transformer [Dwivedi and Bresson, 2020b, Vignac et al., 2023] as the neural network used in GDSS+Transformer [Jo et al., 2022], DiGress [Vignac et al., 2023], and DruM [Jo et al., 2023]. We set the diffusion steps to 1000 for all the diffusion models. For important hyperparameters mentioned in Sec  2.4, we usually set $S_{cale}=0.9$, $S_{hift}=0.09$. and $\eta=[10000,100,30,10]$ for the normalized degrees corresponding to the intervals falling in the interval split by $[1.0,0.8,0.4,0.1]$, respectively. In practice, we set threshold as 0.9 to quantize generated continue adjacency matrix.
A.2 2D molecule generation
Datasets
We utilize two widely-used molecular datasets as benchmarks, as described in Jo et al. [2023]: QM9 [Ramakrishnan et al., 2014], consisting of 133,885 molecules with $N\leq 9$ nodes from 4 different node types and ZINC250k [Irwin et al., 2012], consisting of 249,455 molecules with $N\leq 38$ nodes from 9 node types. Molecules in both datasets have 3 edge types, namely single bond, double bond, and triple bond. Following the standard procedure in the literature [Shi et al., 2020, Luo et al., 2021, Jo et al., 2022, 2023], we kekulize the molecules using the RDKit library [Landrum et al., 2006] and remove the hydrogen atoms from the molecules in the QM9 and ZINC250k datasets.
Evaluation metrics
Following the evaluation setting of  Jo et al. [2022], we generate 10,000 molecules for each dataset and evaluate them with four metrics: the ratio of valid molecules without correction (Val.). Frechet ChemNet Distance (FCD) evaluates the chemical properties of the molecules by measuring the distance between the feature vectors of generated molecules and those in the test set using ChemNet. Neighborhood Subgraph Pairwise Distance Kernel (NSPDK) assesses the quality of the graph structure by measuring the MMD between the generated molecular graphs and the molecular graphs from the test set. Scaffold Similarity (Scaf.) evaluates the ability to generate similar substructures by measuring the cosine similarity of the frequencies of Bemis-Murcko scaffolds [Bemis and Murcko, 1996].
Implementation details
We follow the evaluation setting of  Jo et al. [2022, 2023] to generate 10,000 molecules and evaluate graphs with test data for each dataset. We quote the baselines results from  Jo et al. [2023]. For a fair comparison, we adopt the Graph Transformer [Dwivedi and Bresson, 2020b, Vignac et al., 2023] as the neural network used in GDSS+Transformer [Jo et al., 2022], DiGress [Vignac et al., 2023], and DruM [Jo et al., 2023]. We apply the exponential moving average (EMA) to the parameters while sampling and set the diffusion steps to 1000 for all the diffusion models. For both QM9 and ZINC250k, we encode nodes and edges to one-hot and set $S_{cale}=0.9$, $S_{hift}=0.09$. For $\eta$ modulated in molecule generation, with the help of chemical knowledge, we apply $\eta=[10000,100,100,100,30]$ on carbon-carbon bond, carbon-nitrogen, carbon-oxygen, carbon-fluorine, and other possible bonds, respectively. For $\eta$ of nodes, we apply $\eta=[10000,100,100,30]$ on carbon atom, nitrogen atom, oxygen atom and other possible atoms, respectively. As described in Sec 2.4, applying appropriate $\eta$ for different node types and edge types can prolong the presence of related substructures during the diffusion process. In practice, we set threshold as 0.9 to quantize generated continue adjacency matrix and the value in discrete adjacency matrix is 0 after quantizing if and only if all values in each dimension are all 0.
A.3 Computing resources
For all experiments, we utilized the PyTorch [Paszke et al., 2019] framework to implement GBD and trained the model with NVIDIA GeForce RTX 4090 and RTX A5000 GPUs.
Appendix B Visualization
B.1 Generative process of GBD on general datasets
@@
We visualize generative process of GBD on the Community-small and the Ego-small dataset in Figures 3 and 4, respectively.
@@
B.2 Generated graphs of GBD on 2D molecule datasets

### v2

@@ E.1 Evaluation with larger sample size
As described in Section 4.1, small number of nodes and the insufficient size of the sampled graphs can lead to large standard deviations when evaluating with the reference graph on smaller dataset. Therefore, we attempted to evaluate the large number of generated graphs and report the results in Table 8. We observe that our proposed GBD outperforms previous continuous and discrete diffusion models on both smaller datasets. Furthermore, GBD significantly surpasses the wavelet-based diffusion model (Wave-GD) by a wide margin on the Community-small dataset, as evidenced by both means and standard deviations, Specifically, GBD achieves 85.0%, 90.5%, and 40.0% improvements over Wave-GD in the MMDs means of Degree, Cluster, and Orbit, respectively, indicating that our proposed model is capable of generating smaller graphs that are closer to the data distribution with better stability.
@@
Table 8: Generic graph generation results with enlarged sample (1024 graphs).
@@
E.2 Comparison on various node feature initialization.
Node feature initialization
We initialize node representations using the following node-level features, respectively:
  • Degree (one-hot): Degree with one-hot format is a categorical representation of a node’s degree. It encodes the degree information as a binary vector where each position corresponds to a possible degree value. The position corresponding to the node’s actual degree is set to 1, while all other positions are 0.
  • Degree (normalized): The normalized degree is a continuous representation of a node’s degree, scaled to a value between 0 and 1.
  • Centrality: Here we adopt the normalized betweenness centrality as initial node features and it is a measure of a node’s importance based on its role in connecting other nodes. It quantifies the fraction of shortest paths between all pairs of nodes that pass through the given node. The normalized version scales this value to be between 0 and 1.
  • Eigenvectors: Following Jo et al. (2022), we adopt the two first eigenvectors associated to non zero eigenvalues as initial node features.
@@ Results
As shown in Table 9, GBD outperforms GDSS + TF and ConGress by a large margin in all MMDs when the node representation exhibits sparsity and long-tailedness. Additionally, GBD achieves competitive performance compared with other Gaussian-based diffusion models while the node feature is initializing with Eigenvectors. This demonstrates that our proposed GBD has the ability to model graphs with flexible node features, indicating its potential for modeling graphs with more informative features.
Table 9: The effect of Feature Initialization on Community-small and Ego-small.
Community-samll
Node Feature Degree (one-hot) Degree (normalized) Centralities Eigenvectors
Method Deg. Clus. Orbit. Deg. Clus. Orbit. Deg. Clus. Orbit. Deg. Clus. Orbit.
GDSS+TF 0.008 0.080 0.005 0.009 0.077 0.005 0.010 0.075 0.004 0.005 0.061 0.003
ConGress 0.024 0.072 0.006 0.020 0.076 0.006 0.013 0.079 0.005 0.004 0.067 0.003
GBD 0.002 0.060 0.002 0.004 0.059 0.003 0.003 0.059 0.003 0.004 0.064 0.002
Ego-samll
Node Feature Degree (one-hot) Degree (normlized) Centralities Eigenvectors
Method Deg. Clus. Orbit. Deg. Clus. Orbit. Deg. Clus. Orbit. Deg. Clus. Orbit.
GDSS+TF 0.016 0.020 0.004 0.018 0.023 0.005 0.017 0.026 0.006 0.013 0.015 0.002
ConGress 0.045 0.059 0.015 0.037 0.064 0.017 0.032 0.057 0.014 0.020 0.029 0.011
GBD 0.011 0.014 0.002 0.013 0.017 0.002 0.015 0.012 0.003 0.017 0.016 0.002
Appendix F Visualization
@@
We follow the implementation described in Section 2.3 and the nodes in all adjacency matrices are reordered by decreasing the degree of nodes. Apparently, we can find that edges associated with nodes with large degree will be the first to be identified and then spread in decreasing order of degree on both datasets in Section F.1. It is worth noting that the reverse beta diffusion can converge rapidly, leading to generated graphs with correct topology at an early stage. This shows that our proposed GBD can further explore the potential benefits of beta diffusion, resulting in valid graphs with stability and high quality. For generated molecule graphs shown in Section F.1, we can observe that GBD can successfully generate valid and high-quality 2D molecules, verifying its ability to model attributed graphs. More generated graphs are presented following.
F.1 Generative process of GBD on general datasets
@@
We visualize the generative process of GBD on the Community-small and the Ego-small dataset in Figures 5 and 6, respectively.
@@
F.2 Generative process of GBD on complex graph
We provide the visualization of the complex graph generated by GBD on the Planar and the SBM datasets in Figure 7 and in Figure 8, respectively.
Figure 7: Visualization of the generated graphs of GBD on the Planar dataset.
Figure 8: Visualization of the generated graphs of GBD on the SBM dataset.
F.3 Generated graphs of GBD on 2D molecule datasets